OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1: −5|3x − 2| = −15 Equation #2: |2x − 9| = 11 A.) Equation #1 and Equation #2 have the same number of solutions. B.) Equation #1 has more solutions than equation #2. C.) Equation #1 has fewer solutions than equation #2. D.) None of the statements above describe the number of solutions to equations shown.

OpenStudy (anonymous):

Eq 1: /3x-2/= 3, so x = 5/3 and x = -1/3 Eq 2: 2x - 9 = 11 and 2x - 9 = -11, so x = 10, x = -1 They both have two solutions.